The displacement vectors A and B, when added together, give the resultant vector R, so that R = A + B. Use the data in the drawing and the fact that φ = 21° to find the magnitude R of the resultant vector and the angle θ that it makes with the +x axis.
1
Expert's answer
2020-01-28T13:52:22-0500

"R=\sqrt{6^2+8^2+2(6)(8)\cos{(90+21)\degree}}=8.1\ N"

"8\cos{21}=8.1\cos{\theta}"

"\theta=57.25\degree"